OpenStudy (anonymous):

Solve the equation. Check your answer. 3x – 7 = 5x + 19

OpenStudy (tkhunny):

What is your plan for solving this? Generally, collect 'x' to one side and everything else to the other. Let's see what you get.

OpenStudy (anonymous):

3x-7=5x+19 now get all the "x" on one side 5x-3x=-7-19 2x=-26 x=-13

OpenStudy (tkhunny):

Sweet! Are you doubting? Go check it out inthe ORIGINAL problem statement and you'll be done.

OpenStudy (anonymous):

is it right

OpenStudy (tkhunny):

You tell me. That's why you check it in the original problem statement. If it doesn't work, somehting went wrong along the way. Since checking is in the problem statement, I must first encourage you to do it. Substitute on both sides and simplfy to see if you get the same thing on both sides.

OpenStudy (anonymous):

okmg can somebody tell me if it's right

OpenStudy (anonymous):

omg

OpenStudy (tkhunny):

It can't be right until you complete the assignment! I don't understand why you won't do it. 3x – 7 = 5x + 19 x = -13 3(-13) – 7 = 5(-13) + 19 Check it out! If it works, you are done and you can tell yourself if it is correct. Don't you think that would be of some value?